Find the natural frequencies of the system shown on the right, with m1=m, m2=2m, k1=k, and k2=2k. Determine the response of the system when k=1000 N/m, m=20 kg, and the initial values of the displacements of the masses m1 and m2 are l and -1 respectively.


1
Expert's answer
2021-05-18T16:34:01-0400

The equation of motion for m1

"m_1\ddot{x} _1+(k_1+k_2)x_1-k_2x_2=0"

Substituting "- \omega^2x_1" for "\ddot{x} _1 \implies -\omega^2m_1x_1+(k_1+k_2)x_1-k_2x_2=0"

The equation of motion for m2

"m_2\ddot{x} _2+(k_1+k_2)x_2-k_2x_2=0"

Equation 1 and 2 in form of a matrix

"\begin{bmatrix}
 -\omega^2m_1+k_1+k_2 & -k_2 \\
 -k_2 & k_2-m_2\omega^2
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 
\end{bmatrix}= \begin{bmatrix}
 0 \\
 0 
\end{bmatrix}"

"\omega^4-(\frac{k_1+k_2}{m_1}+\frac{k_2}{m_2})\omega^2+\frac{k_1k_2}{m_1m_2}=0"

"\omega^2_1,\omega^2_2 = \frac{k_1+k_2}{2m_1}+\frac{k_2}{2m_2} \mp \sqrt{\frac{1}{4}(\frac{k_1+k_2}{m_1}+\frac{k_2}{2m_2})^2-\frac{k_1k_2}{m_1m_2}}"

Substituting in the values

"\omega^2_1,\omega^2_2 = \frac{1000+2000}{2*20}+\frac{2000}{2*40} \mp \sqrt{\frac{1}{4}(\frac{1000+2000}{20}+\frac{2000}{2*20})^2-\frac{1000*2000}{20*40}}"

"\omega_1=3.6606rad/s,\omega_2 =13.6603 rad/s"

"r_1=\frac{X_2}{X_1}=\frac{-m_1\omega_2^2+k_1+k_2}{k_2}=\frac{-20*13.6603^2+1000+2000}{2000}=-0.366"

"r_2=\frac{X_2}{X_1}=\frac{-m_1\omega_2^2+k_1+k_2}{k_2}=\frac{-20*13.6603^2+1000+2000}{2000}=-0.366"

"x_1=-0.366 cos 3.6603t-1.366cos13.6603t"

"x_2=-0.5 cos 3.6603t+0.5cos13.6603t"